Circular No. 8000

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2002 U1 (SOHO) Further to IAUC 7991, D. Hammer reports measurements for a Kreutz

sungrazing comet found by X. Leprette on C2 coronagraph images at the SOHO website. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-U08.

2002 UT  α_{2000}  δ_{2000}
Oct. 16.329  13^{h}16.^{m}3  − 9^{o}36′

COMET C/2002 U2 (LINEAR) M. Bezpalko, Lincoln Laboratory, Massachusetts Institute of Technol-

ogy, reports the LINEAR discovery of a comet with a tail in p.a. ∼ 270^{o} (discovery observation below). L. Šarounová, Ondřejov, reports that images taken with the 0.65-m reflector on Oct. 26.0 UT show a 10′′ coma (m_{1} = 15.5) and a narrow tail at least 2′ long in p.a. 280^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Oct. 25.43912  8^{h}09^{m}45.^{s}15  +72^{o}37′35′′.8  17.3

Additional astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2002-U33.

T = 2003 Jan. 3.945 TT  ω = 100.^{o}305
Ω = 39.082 }2000.0
q = 1.15645 AU  i = 59.455

SUPERNOVA 2002ha IN NGC 6962 W. M. Wood-Vasey, Lawrence Berkeley National Laboratory, reports

that NEAT images do not show SN 2002ha (cf. IAUC 7997) on Oct. 18.3 UT (limiting mag 19.85 at S/N = 3, after subtraction of galaxy light from images taken on Sept. 20).

V4742 SAGITTARII Visual magnitude estimates: Sept. 23.014 UT, 9.2 (J. Garcia, Rama

Caida, Argentina); 29.924, 10.4 (R. Y. Shida, São Paulo, Brazil); Oct. 9.503, 11.5 (A. Pearce, Nedlands, W. Australia); 14.510, 11.7 (Pearce).

2002 October 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green